# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v1

@@ Abstract
Advancements in image segmentation play an integral role within the greater scope of Deep Learning-based computer vision. Furthermore, their widespread applicability in critical real-world tasks has given rise to challenges related to the reliability of such algorithms. Hence, uncertainty quantification has been extensively studied within this context, enabling expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision making. Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field. This work provides a comprehensive overview of probabilistic segmentation by discussing fundamental concepts in uncertainty that govern advancements in the field as well as the application to various tasks. We identify that quantifying aleatoric and epistemic uncertainty approximates Bayesian inference w.r.t. to either latent variables or model parameters, respectively. Moreover, literature on both uncertainties trace back to four key applications; (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) active learning. Then, a discussion follows that includes an overview of utilized datasets for each of the applications and comparison of the available methods. We also highlight challenges related to architectures, uncertainty-based active learning, standardization and benchmarking, and recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
@@
image segmentation, uncertainty quantification, probability theory
@@ I Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1]. With the rapid development of Convolution Neural Networks (CNNs), Deep-Learning based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7]. Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions. This ignorance diminishes the reliability and interpretability of such models. For example, difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
@@
Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty. Furthermore, explicitly modeling the output likelihood distribution enables expressing the aleatoric uncertainty. Notably, literature mention that determining the nature of uncertainty is not often straightforward. For example, Hüllermeier et al. [12] mentioned that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”. This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “In one model an addressed uncertainty may be aleatory, in another model it may be epistemic”. Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
@@
The merits of uncertainty quantification have fortunately been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview. Nonetheless, most surveys take perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19]. There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications. Furthermore, the abundance of available works can often be overwhelming for both new-coming and seasoned researchers. This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are clarified through standardized notation. After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks. Additionally, they will have developed comprehensive understanding of challenges and unexplored avenues in the field.
@@
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section II, where closely related surveys are also referenced. Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III. The role of these concepts in image segmentation will be treated in Section IV and V, which includes all architectures and approaches with significant impact on the field. We then consider the applications that use these uncertainty estimates in Section VI. Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII and we conclude in Section VIII. For a brief overview of the sections, refer to Figure 1.
@@
Figure 1: Overview of the sections.
@@ II Background
As summarized by Minaee et al. [20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watersheds [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29]. After the application of CNNs [30], the domain of image segmentation underwent rapid developments. Notably, the Fully Convolutional Network (FCN) [3] adapted the AlexNet [31], VGG16 [32] and GoogLeNet [33] architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 [34], and the MobileNetv3 [35] have also been commonly used.
As the research progressed, increasing success has been observed with encoder-decoder models [36, 4, 37, 2]. Initially developed for the medical applications, Ronneberger et al. [2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures. In fact, reports in recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
Semantic Segmentation focuses solely on assigning a class label(s) to each pixel and is particularly suitable for amorphous or uncountable subjects of interest. In contrast, Instance Segmentation not only detects, but also delineates individual objects within the image. This form of segmentation is more applicable when identifying and outlining countable instances of objects. The third category, Panoptic Segmentation, combines both class and instance level classification [20].
@@ III Probabilistic Image Segmentation
Let random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, where $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively.
@@ III-A Bayesian inference
Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
[…]
As evident in Equation (2), both the variability in the empirical data and the inferred parameters of the model influence the variance of the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy defined as
$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\mathbb{E}\,\left[-\log p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\right]$ (3)
or variance
$\mathrm{Var[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]}=\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})^{2}\,]-\mathbb{E}\,[\,p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})\,]^{2}.$ (4)
@@ III-B Conventional segmentation
The so-called “deterministic” segmentation networks are trained by Maximum Likelihood Estimation (MLE) as
$\bm{\theta}_{MLE}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (5)
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the variance of the model parameters approaches zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution. This is typically done through Bayesian Maximum A Posteriori (MAP) estimation with
$\bm{\theta}_{MAP}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (6)
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the L2-norm (also known as ridge regression or weight decay) or L1-norm of $\bm{\theta}$, respectively [41, 42]. Additionally, the output of such models can be interpreted as a probability distributiont For instance, let a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$, where we denote the input image and segmentation dimensionality as $\mathbb{R}^{C\times D}$ and $\mathbb{R}^{K\times D}$, respectively, for concise notation. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
@@
with channel-wise indexing over in the denominator, which is commonly known as the SoftMax activation. See Figure 2 for an illustration. While not referred to as such in common nomenclature, this is probabilistic modeling in the technical sense, and the approximated distribution can represent and localize uncertain regions. However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks. This challenge can be addressed either from the aleatoric or epistemic perspective, inferring Bayes theorem w.r.t. the hidden latent variables or model parameters, respectively. Both approaches have relevant specific use-cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
@@
Figure 3: Interpretation of the likelihood function in segmentation models. Color intensity in (a) reflect the normalized confidence values that can be interpreted as probabilities. Maximum likelihood thresholding can be applied to obtain (b). However, the coherence of the segmentation suffers when one attempts to sample, as is shown in (c).
@@ IV Aleatoric uncertainty
with Dirac-delta function $\delta$, and arbitrary function $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$, and is inherently probabilistic due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, detecting substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition. The possible input-dependency drives further categorization in to either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, one can model a distribution parameterized by the output of a CNN (Section IV-A). Also, generative models have extensively been used, where the data likelihood is learned through latent variables (Section IV-B). Finally, image augmentation during test-time inference can also be applied to obtain a notion of aleatoric uncertainty (Section IV-C).
@@ IV-A Pixel-level sampling
A valid approach employs direct pixel-level uncertainty in the annotations. As discussed before, uncertainty can be quantified in case the independence assumption holds. In this case, it is important to ensure proper calibration, which is discussed in Section IV-A1. In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
@@ IV-A1 Independence
The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration. Ideally, the empirical accuracy of a model should approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over- confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Different methods can be used to measure calibration, but often introduce their own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy and confidence bins as
$ECE=\sum_{b=1}^{B}\frac{n_{b}}{N}|\operatorname{acc}(b)-\operatorname{conf}(b)|,$ (10)
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under- confidence. Furthermore, Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
@@
Figure 4: Reliability diagram indicating if the model confidence accurately reflect the empirical probabilities.
Contemporary SoftMax-activated neural networks often portray a significantly incongruous reflection of the true data uncertainty because of negative log-likelihood optimization and techniques such as batch normalization, weight decay and other forms of regularization [9]. Most calibration techniques are post-hoc, i.e. they occur after training and thus require a separate validation set. For example, Temperature Scaling [9] has been applied in a pixel-wise manner for segmentation problems [43]. Nonetheless, some methods, such as Label Smoothing [44, 45] or using Focal loss [46] can directly be applied on the training data. Furthermore, overfitting has often been considered to be the cause of overconfidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].

### v3

@@ Abstract
Advancements in image segmentation play an integral role within the broad scope of Deep Learning-based Computer Vision. Furthermore, their widespread applicability in critical real-world tasks has resulted in challenges related to the reliability of such algorithms. Hence, uncertainty quantification has been extensively studied within this context, enabling the expression of model ignorance (epistemic uncertainty) or data ambiguity (aleatoric uncertainty) to prevent uninformed decision-making. Due to the rapid adoption of Convolutional Neural Network (CNN)-based segmentation models in high-stake applications, a substantial body of research has been published on this very topic, causing its swift expansion into a distinct field. This work provides a comprehensive overview of probabilistic segmentation, by discussing fundamental concepts of uncertainty quantification, governing advancements in the field as well as the application to various tasks. Moreover, literature on both types of uncertainties trace back to four key applications: (1) to quantify statistical inconsistencies in the annotation process due ambiguous images, (2) correlating prediction error with uncertainty, (3) expanding the model hypothesis space for better generalization, and (4) Active Learning. An extensive discussion follows that includes an overview of utilized datasets for each of the applications and evaluation of the available methods. We also highlight challenges related to architectures, uncertainty quantification methods, standardization and benchmarking, and finally end with recommendations for future work such as methods based on single forward passes and models that appropriately leverage volumetric data.
@@
image segmentation, uncertainty quantification, probabilistic modeling, epistemic, aleatoric
@@ I Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest [1]. With the rapid development of Convolution Neural Networks (CNNs), Deep Learning-based image segmentation has seen major advancements and gained significant interest over time [2, 3, 4], obtaining impressive scores with large-scale segmentation datasets [5, 6, 7]. Nevertheless, such methodologies rely on extensive assumptions and relaxations on the Bayesian learning paradigm, omitting crucial information on the uncertainty associated with the model predictions. This omission diminishes the reliability and interpretability of such models. For example, a difficult distinction between classes in real-time automotive scenarios can result in disastrous consequences or uncertain lesion malignancy prediction may significantly impact the decision-making of invasive treatments.
@@
Extensive efforts have been made to align modern neural network optimization with Bayesian Machine Learning [8, 9, 10, 11], such as learning parameter densities, rather than point estimates, to include a notion of epistemic uncertainty. Furthermore, explicitly modeling the output likelihood distribution enables expressing the aleatoric uncertainty. Notably, literature mention that determining the nature of uncertainty is not often straightforward. For example, Hüllermeier et al. [12] stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred”. This sentiment is also shared by Kiureghian and Ditlevsen [13], noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic”. Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic.
@@
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation and underscore the importance of a rigorous literature overview. Nonetheless, most surveys take the perspective from the medical domain [14, 15], often for specific modalities [16, 17, 18, 19]. There is a notable absence of a comprehensive overview in this field, relating theoretical foundation to the multitude of applications. Furthermore, the abundance of available works can be often overwhelming for both emerging and seasoned researchers. This study seeks to fill this gap in literature and contributes to the subject area by providing a curated overview, where various concepts are discussed through standardized notation and clarified by mutual comparison. After reading this work, readers will be able to discern various forms of uncertainty with their pertinent applications to segmentation tasks. Additionally, they will have developed comprehensive understanding of challenges and unexplored avenues in the field.
@@
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section II, where closely related surveys are also referenced. Then, the theoretical framework and notation that governs the remainder of the paper is introduced in Section III. The role of these concepts in image segmentation will be treated in Sections IV and V, which include all architectures and approaches with significant impact on the field. We then consider the applications that use these uncertainty estimates in Section VI. Finally, this overview will be further discussed together with key challenges and future recommendations in Section VII, which is concluded in Section VIII. Figure 1 illustrates a brief overview of the sections.
@@
Figure 1: Overview of the sections. The leafs of the presented hierarchical tree are related to subsections of the article.
@@ II Background
This section highlights a selection of literature crucial to the development of current probabilistic segmentation models. While literature on segmentation itself is vast, this decision has resulted in a limited collection of studies.
As summarized by Minaee et al. [20], semantic segmentation has been performed using methods such as thresholding [21], histogram-based bundling, region-growing [22], k-means clustering [23], watershedding [24], to more advanced algorithms such as active contours [25], graph cuts [26], conditional and Markov random fields [27], and sparsity-based methods [28, 29]. After the application of CNNs [30], the domain of image segmentation underwent rapid developments. Notably, the Fully Convolutional Network (FCN) [3] adapted the AlexNet [31], VGG16 [32] and GoogLeNet [33] architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 [34], and the MobileNetv3 [35] have also been commonly used.
As the research progressed, increasing success has been observed with encoder-decoder models [36, 4, 37, 2]. Initially developed for the medical applications, Ronneberger et al. [2] introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for segmentation architectures, particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net and nnU-Net [38] still outperform more contemporary and complex models [39, 40].
Semantic segmentation focuses solely on assigning class labels to each pixel and is particularly suitable for amorphous or uncountable subjects of interest. In contrast, instance segmentation not only detects, but also delineates individual objects within the image. This form of segmentation is more applicable when identifying and outlining countable instances of objects. The third category, panoptic segmentation, combines both class and instance-level classification [20].
@@ III Probabilistic Image Segmentation
This section describes the fundamentals of Bayesian inference and its relationship with conventional models for segmentation. The provided details introduce standard variables and notations as we proceed throughout the manuscript.
@@ III-A Bayesian inference
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{R}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then $\mathbf{Y}$ can be considered as the ground-truth of a $K$-class segmentation task and $\mathbf{X}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
$p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})=\int\underbrace{p(\mathbf{Y}|\mathbf{x}^{*},\bm{\theta})}_{\text{Data}}\overbrace{p(\bm{\theta}|\mathcal{D})}^{\text{Model}}\mathrm{d}\bm{\theta}.$ (2)
As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. A straightforward approach to quantify any of these uncertainties is achieved through obtaining the predictive entropy, which can similarly be decomposed as
$H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]=\overbrace{I[\mathbf{y},\bm{\theta}|\mathbf{x}^{*},\mathcal{D}]}^{\text{epistemic}}+\underbrace{\mathbb{E}_{q(\bm{\theta}|\mathcal{D})}[\,H[\mathbf{y}|\mathbf{x}^{*},\bm{\theta}]\,]}_{\text{aleatoric}}.$ (3)
@@ III-B Conventional segmentation
Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks, which are trained by Maximum Likelihood Estimation (MLE) and is specified as
$\bm{\theta}_{\text{MLE}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta}),$ (4)
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution. This is typically achieved through Bayesian Maximum A Posteriori (MAP) estimation with
$\bm{\theta}_{\text{MAP}}=\operatorname*{arg\,max}_{\bm{\theta}}\log p(\mathbf{y}|\mathbf{x},\bm{\theta})+\log p(\bm{\theta}).$ (5)
For example, assuming Gaussian or Laplacian priors can be a consequence of regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively [41, 42]. Additionally, the output of such models can be interpreted as a probability distribution. For instance, consider spatial image dimensions $D$ and a CNN model $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ with parameters $\bm{\theta}$, such that $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, the output of such model can be regarded as parameters of a Probability Mass Function (PMF) through
@@
with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2). This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature. In fact, the approximated distribution can represent and localize uncertain regions.
However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. At the very core, this limitation (See Figure 3) has driven research on probabilistic segmentation, enabling the sampling of spatially coherent segmentation masks. This challenge can be addressed either from the aleatoric or epistemic perspective. Both approaches have relevant specific use cases, and distinct merits and drawbacks, as will become clear in the upcoming sections.
@@
Figure 3: Illustration of the likelihood function in segmentation models. Color intensities reflect the normalized confidence values that can be interpreted as probabilities. (a) The continuous output results in a horizontal gradient. (b) Maximum likelihood thresholding can be applied. (c) However, the coherence of the segmentation suffers when sampling.
@@ IV Aleatoric uncertainty
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty.
In most practical scenarios, aleatoric uncertainty quantification methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN (Section IV-A). Also, generative models have been used extensively, where the data likelihood is learned through latent variables (Section IV-B). Finally, image augmentation during test-time inference can be also applied to obtain a notion of aleatoric uncertainty (Section IV-C).
@@ IV-A Pixel-level sampling
A valid approach employs direct pixel-level uncertainty in the annotations. In case the independence assumption holds, it is important to ensure proper calibration of the model, which is discussed in Section IV-A1. In another method, the spatial correlation is explicitly modeled across the pixels of the segmentation mask and results in spatially coherent samples, which will be treated in Section IV-A2.
@@ IV-A1 Independence
The normalized confidence values that result from SoftMax activation can only be interpreted as probabilities after proper validation, which is referred to as model calibration. Ideally, the empirical accuracy of a model should be approximately equal to the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Different methods can be used to measure calibration, but often introduce their own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
$E_{\text{ECE}}=\sum_{b=1}^{B}\frac{n_{b}}{N}|\text{acc}(b)-\text{conf}(b)|,$ (9)
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence. Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.
@@
Figure 4: Visualization of a reliability diagram indicating if the model confidences accurately reflect the empirical probabilities.
Contemporary SoftMax-activated neural networks often portray a significantly incongruous reflection of the true data uncertainty because of negative log-likelihood optimization and techniques such as batch normalization, weight decay and other forms of regularization [9]. Most calibration techniques are post-hoc, i.e. they occur after training and thus require a separate validation set. For example, Temperature Scaling [9] has been applied in a pixel-wise manner for segmentation problems [43]. Nonetheless, some methods, such as Label Smoothing [44, 45] or using the Focal Loss [46] can be directly applied on the training data. Furthermore, over-fitting has often been considered to be the cause of over-confidence [47, 48] and erroneous pixels can therefore be penalized through regularizing low-entropy outputs [49].